A chemist needs 2 liters of 14% solution. He has 10% solution and a 20% solution to form this mixture. How much of each should he use to get the mixture?
I have already found that he needs 0.8 liters of the 20% solution, I can't seem to get the first one. Thank you!

Answer by rothauserc(4718)

x + y = 2
.10x +.20y = 2*.14 = .28
solve fist equation for x and substitute in second equation
.10*(2 -y) +.20y = .28
.20 -.10y + .2oy = .28
.10y = .08
y = .8 and x = 1.2
so we need .8 liters of the 20% solution and 1.2 liters of the 10% solution
check the answers by substituting in the second equation
.10*1.2 +.20*.8 = .28
.12 + .16 = .28
.28 = .28
answers check